You have a 0.025 L solution of piperazine (perhydro-1,4-diazine), a weak dibasic base. The K, values for its conjugate acid a Dont need (a) or (b) thank you
Answer #1

pKb1 = -Log(4.65*10-6) = 5.33, pKb2 = -Log(1.86*10-10) = 9.73

The concentration of salt at first equivalence point = 0.0025 mol/(0.025 + 0.05416) L = 0.0316 M

The concentration of salt at second equivalence point = 0.0025 mol/(0.025 + 2*0.05416) L = 0.0188 M

Part c) Moles of piperazine in the initial solution = 0.025 L * 0.1 mol/L = 0.0025 mol

Part d) Concentration of intitial solution = 0.1 M

Part e) pH of initial solution = 14 - 1/2 (pKb1 - Log0.1) = 14 - 1/2 (5.33 + 1) = 10.83

Part f) pH at first equivalence point = 7 - 1/2 (pKb1 + Log[salt]) = 7 - 1/2 (5.33 + Log0.0316) = 5.08

Part g) pH = 1/2 (14-5.33 + 14-9.73) = 6.47

Part h) pH at the second equivalence point = 7 - 1/2 (pKb2 + Log[salt]) = 7 - 1/2 (9.73 + Log0.0188) = 3